OpenStudy (anonymous):

-a/10-8=-24 two step equation

OpenStudy (anonymous):

\[\frac{-a}{10}-8=-24\] \[\frac{-a}{10}=-24+8\] \[\frac{-a}{10}=-16\] \[-a=10\times -16\] \[-a=-160\] \[a=160\]